A THEORETICAL PREDICTION OF ELECTRONIC TRANSITIONS IN C_{3}

\underline{DAVID W. SCHWENKE}, GALINA M. CHABAN, NASA Ames Research Center, Moffett Field, CA 94035.

We have carried out first principle calculations of electronic transitions in C_{3}. Low lying electronic states have linear or near linear minima, and thus are characterized by D_{∞v} symmetry labels. The ground state has ^{1}Σ_{g}^{+} symmetry, and can be resonably treated by triatomic bound state codes that assume the Born-Oppenheimer approximation is valid . The lowest excited state has ^{1}Π_{u} symmetry, and explicit coupling between electronic and nuclear angular momentum must be considered: the Renner-Teller effect. We treat this coupling exactly so our results are independent of the choice of coordinates. Furthermore, the ^{1}Π_{u} state is perturbed by the lowest triplet state, and we include this coupling as well.